A particular group of men have heights with a mean of 177 cm and a standard deviation of 7 cm. JohnJohn had a height of 193 cm. a.a. What is the positive difference between JohnJohn​'s height and the​ mean? b.b. How many standard deviations is that​ [the difference found in part​ (a)]? c.c. Convert JohnJohn​'s height to a z score. d.d. If we consider​ "usual" heights to be those that convert to z scores between minus−2 and​ 2, is JohnJohn​'s height usual or​ unusual?
Answer #1

Ans:

a)

Positive difference=193-177=16

b)

k=16/7=2.29

2.29 standard deviations

c)

z score=(193-177)/7=2.29

d)unusual,as z score is more than 2.